# Integrating Neural Operators with Diffusion Models Improves Spectral Representation in Turbulence Modeling

### v1

@@ 3 Discussion
The energy spectra in Figure 2 to Figure 7 demonstrate that the diffusion model conditioned on the neural operator performs better than the neural operator alone. However, these plots are constructed for each time step separately and do not contain information about the cross-correlation of the states across time steps. We address this question by performing proper orthogonal decomposition (POD) on 10 timesteps predicted by the neural operator, as well as the diffusion model conditioned on the neural operator, and compare it with the corresponding state of the ground truth. Figure 8 illustrates the decay of the eigenvalues and compares the POD modes of the true data with the other two approaches for the Schlieren velocimetry case. The proposed framework of the diffusion models conditioned on the neural operator has better aligned eigen spectrum and eigen modes with the ground truth than the neural operators alone. The POD analysis of Kolmogorov flow (Figure A.1), buoyancy-driven flow (Figure A.2), turbulent airfoil (Figure A.3) and 3D turbulent jet (Figure A.4) are also provided in Appendix A.
Fig.  8: POD analysis of Schlieren velocimetry. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator and diffusion model conditioned on neural operator (MATCHO) with the true solution obtained from Schlieren velocimetry. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
@@
Summarizing, we propose a framework that utilizes neural operators as effective priors for diffusion models. Our results demonstrate that the proposed framework that blends neural operators with diffusion models significantly improves the energy spectrum compared to the neural operators alone and has been tested across different types of neural operator architectures - FNO, U-Net, and MATCHO, over a variety of test cases. The framework utilized here is general, and therefore, any advances in the quality, computational efficiency, or scalability of either the neural operator or the diffusion models in the future can be applied seamlessly and could potentially improve the proposed framework.
@@ 4.1 Neural Operators
Neural Operators are a class of architectures in deep learning, that maps between infinite dimensional functional spaces. Consider an operator $\mathcal{G}$, that maps from the input function $X$ to the output function $Y$, i.e., $\mathcal{G}:X\rightarrow Y$. Then any machine learning architecture that is used to approximate this functional mapping to predict $y\in Y$ for a given $x\in X$ can be defined as a neural operator. Neural operator is a popular framework for encoding dynamics and making predictions under some given conditions. But like any other neural network architectures, they suffer from spectral bias.
Lemma 4.1. Let $F:X\rightarrow Y$ be a functional mapping with $X,Y\in H^{s}(\Omega)$, where $H^{s}(\Omega)$ is the Sobelov space of functions defined on $\Omega$. Let $\mathcal{G}(\theta_{NO})$ be a neural operator parameterized by $\theta_{NO}$, that is approximating the functional space $F$. Then $\mathcal{G}$ learns the lower frequency components of the functional space $F$ more effectively.
Sketch of Proof: We want to approximate the functional space $F(X)=Y$, then if we take the Fourier transform for this function, we have
@@
where $\hat{F}(\xi)$ represents the amplitude of the the frequency component $\xi$. Now let $\mathcal{G}(\theta_{NO},X)$ is the operator approximating $F(X)$, typically the loss function $\mathcal{L}(\theta_{NO})$ used for optimizing such networks is given by
$\mathcal{L}(\theta_{NO})=\frac{1}{N}\sum_{n=1}^{N}(F(x_{n})-\mathcal{G}(\theta_{NO};x_{n}))^{2},$ (7)
where $x_{n}\in X$, and $n=1,2,\cdot\cdot\cdot,N$ are the discrete points where the functional evaluation are available for training. We can also express $\mathcal{G}$ in terms of its discrete Fourier components as
$\hat{\mathcal{G}}(\theta_{NO};\xi)=\frac{1}{N}\sum_{n=1}^{N-1}\mathcal{G}(\theta_{NO};x_{n})e^{-2\pi i\xi n/N},$ (8)
@@
$\mathcal{G}(\theta_{NO};x_{n})=\frac{1}{N}\sum_{n=1}^{N-1}\hat{\mathcal{G}}(\theta_{NO};\xi)e^{2\pi i\xi n/N}.$ (9)
@@
$\nabla_{\theta_{NO}}\mathcal{L}(\theta_{NO})=\frac{-2}{N}\sum_{n=1}^{N}(F(x_{n})-\mathcal{G}(\theta_{NO};x_{n}))\nabla_{\theta_{NO}}\mathcal{G}(\theta_{NO};x_{n}).$ (11)
@@
$\frac{d}{dm}(\hat{\mathcal{G}}(\theta_{NO};\xi))\propto-\alpha\hat{\nabla}_{\theta_{NO}}\mathcal{L}(\theta_{NO}).$ (12)
$\alpha\hat{\nabla}_{\theta_{NO}}\mathcal{L}(\theta_{NO})$ is larger for smaller $\xi$. Therefore, the operator $\mathcal{G}$ learns the low frequency components of the functional space $F(X)$ more effectively compared to higher frequency components, which is known as the spectral bias. Hence like any other networks, operators suffer from spectral bias, which leads to ineffective refinement of the higher frequency component, leading, in turn, to sharp decay of the power spectrum associated with the output predicted by any neural operator. Although neural operators fail to capture the high frequency, it captures effectively the mean flow dominated by the low wave numbers. This acts as an ideal prior for diffusion models to mimic the true flow consisting of all the frequencies.
@@ 4.2 Score-based diffusion models
The overall goal of a score ($s$) based diffusion model is to estimate the score function defined by $s_{\theta_{D}}(\mathbf{X})=\nabla_{X}\log p(\mathbf{X})$, where $\theta_{D}$ are the parameters of the diffusion model and $p$ is the probability density of $\mathbf{X}$. Since the data distribution is not known explicitly, and there is a chance that the data might lie on a lower dimensional manifold, the score function may be ill defined in regions where there is no data. To circumvent this issue, perturbation of the data is done by adding Gaussian noise, which ensures that the score is well defined across the entire space by smoothing the distribution. Hence, the score function gives us a direction to move towards the region of higher probability. However the direct way to sample from the distribution is missing here. In order to fix this problem, Langevin Dynamics was used in [54]. Langevin Dynamics ensures the convergence of the generated samples to the true underlying distribution. It balances the inherent dynamics between deterministic movement, which is driven by the gradient of log of the probability, with random exploration given by the noise levels. In our proposed method, we use the score function, which is additionally conditioned on the output of the neural operator $\mathbf{Y}=\mathbf{\mathcal{G}}(\theta_{NO},\mathbf{X})$ given by
$s_{\theta_{D}}(\mathbf{X},\sigma,\mathbf{Y})=\nabla_{X}\log p(\mathbf{X}|\mathbf{Y},\sigma),$ (13)
@@
where $\sigma$ is the noise level. The modified score function now considers the joint distribution of $\mathbf{X}$ and $\mathbf{Y}$. This ensures that the the generated samples from the diffusion model are consistent with both the structures defined by $\mathbf{Y}$ and the data distribution that is learned by the diffusion model. The update rule for Langevin dynamics can be shown as
$\mathbf{X}_{i+1}=\mathbf{X}_{i}+\frac{\varepsilon}{2}s_{\theta_{D}}(\mathbf{X}_{i},\sigma_{t},\mathbf{Y})+\sqrt{\varepsilon}z_{i},$ (14)
@@
where $\varepsilon$ is the step size and $z_{i}$ is the noise component. From Eqns 13 and 14, we can see that as $\sigma_{i}$ decreases, the score function $s_{\theta_{D}}$ focuses on recovering the high-frequency details which was smoothen out by the noise. In the frequency domain Eqn 14 can be shown as
$\hat{\mathbf{X}}_{i+1}=\hat{\mathbf{X}}_{i}+\frac{\varepsilon}{2}\hat{s}_{\theta_{D}}(\mathbf{X}_{i},\sigma_{i},\mathbf{Y})+\sqrt{\varepsilon}\hat{z}_{i},$ (15)
@@
$u(\mathbf{X},t)=\sum_{n=1}^{\infty}a_{n}(t)\phi_{n}(\mathbf{X}),$ (16)
where $\phi_{n}(\mathbf{X})$ are the basis functions and $a_{n}(t)$ are the time-dependent coefficients. Let $U_{NO}(\mathbf{X},t)$ and $U_{D}(\mathbf{X},t)$ represents the output of the neural operator and the diffusion model respectively, such that
$U_{NO}(\mathbf{X},t)+U_{D}(\mathbf{X},t)\simeq\sum_{n=1}^{\infty}a_{n}(t)\phi_{n}(\mathbf{X}).$ (17)
If $k_{\xi}$ represents the wave number which determines the basis function approximated by $U_{NO}$ such that
$E_{U_{NO}}(k_{i})=\frac{1}{N}\left|\sum_{(k_{n}\leq k_{\xi})}a_{n}(t)\hat{\phi_{n}}(k_{i})\right|^{2},$ (18)
where $E_{U_{NO}}(k_{i})$ is the energy spectrum related to $U_{NO}$ in the discrete Fourier domain at wave number $k_{i}$, then $\left|u(\mathbf{X},t)-(U_{NO}(\mathbf{X},t)+U_{D}(\mathbf{X},t))\right|\rightarrow 0$ as $k_{\xi}$ increases.
@@
$U_{NO}(\mathbf{X},t)+U_{D}(\mathbf{X},t)$ $\simeq\sum_{n=1}^{\infty}a_{n}(t)\phi_{n}(\mathbf{X})$ (19)
$\implies U_{NO}(\mathbf{X},t)+U_{D}(\mathbf{X},t)$ $=\sum_{n:k_{n}\leq k_{\xi}}a_{n}(t)\phi_{n}(\mathbf{X})+\sum_{n:k_{n}>k_{\xi}}a_{n}(t)\phi_{n}(\mathbf{X}).$ (20)
In the discrete Fourier space we can write it in terms of the energy spectrum as
$E_{U_{NO}}(k_{i})+E_{U_{D}}(k_{i})=\frac{1}{N}\left|\sum_{(k_{n}\leq k_{\xi})}a_{n}(t)\hat{\phi_{n}}(k_{i})\right|^{2}+\frac{1}{N}\left|\sum_{(k_{n}>k_{\xi})}a_{n}(t)\hat{\phi_{n}}(k_{i})\right|^{2}.$ (21)

### v2

@@ 3 Discussion
Impact of Neural Operator Accuracy on Diffusion Model
A careful analysis of the buoyancy-driven flow results in Figure 4 reveals a discrepancy between the true and estimated states, especially at later time steps, due to the inability of the neural operator to estimate the high-frequency content that has low energy. At each autoregressive step this error cascades up the energy spectrum towards the low-frequency content that bear higher energy. Although the diffusion model offers a better-aligned energy spectrum by recovering the missing frequencies, the error introduced by the neural operator does not go away. This motivated us to investigate the impact of neural operator accuracy on the diffusion model by varying the number of trainable parameters of neural operator as 2M, 8M, 32M, and 128M. The diffusion model has only 2M parameters and is not varied. We analyze the concentration field $d$ at $t=12$, predicted by all the neural operators and diffusion models in Figure 8. Increasing the number of trainable parameters enables the neural operator to better learn the high frequency content, subsequently leading to lower errors. An improved neural operator also serves as a better prior to the diffusion model. The relative MSE of $d$ predicted by the neural operator alone and by the diffusion model conditioned on the neural operator is reduced by increasing the number of trainable parameters in the neural operator. However, the diffusion model’s predictions have slightly higher mean errors than the neural operator’s output. Although the diffusion model helps to recover the missing frequencies, they do not occur at all the correct locations resulting in slightly higher errors on the average, as shown in the centerline plots due to misalignment or phase errors. Regarding the energy spectrum, the diffusion model offers an order-of-magnitude improvement over the neural operator on the relative MSE. To summarize, both the field and spectrum error of the diffusion model’s predictions decreases if we have an improved prior from a more accurate neural operator. If there is a cost or memory constraint, for example, 4M trainable parameters, the diffusion model (2M) conditioned on the neural operator (2M) can still offer predictions with well-aligned energy spectrum.
Fig.  8: Sensitivity to neural operator size. We compare the states at $t=12$ for a sample from test dataset predicted by diffusion model conditioned on neural operators with 2M, 8M, 32M, and 128M parameters. The energy spectrum and centreline plots are also provided. b) and c) represents the change in relative MSE of the concentration field $d$ and relative MSE of the log spectrum of $d$.
Is field MSE the perfect evaluation metric?
We have added Table 1 in order to address this question. We show two metrics - the field error and the energy spectrum error. The field error corresponds to the relative MSE between the ground truth and the prediction, and the energy spectrum error corresponds to the relative MSE between the log power spectrum of the ground truth and the prediction. The field error of NO (Neural Operator) is slightly better than NO+DM (Neural Operator + Diffusion Model). The energy spectrum error of NO+DM is better than NO. In order to understand this, we need to carefully examine what each metric represents. The field error does not account for all frequencies uniformly. It is more sensitive to those frequencies that carry higher energy, and is less impacted by the frequencies that bear lower energy. In all the turbulent fluid flow problems considered in this study, the energy spectrum reveals that most of the energy is carried by the lower frequencies. Therefore, the field error metric dominantly represents the ability of the surrogate model to accurately reconstruct the mean flow represented by the lower frequencies. The high frequency small-scale structures that bear less energy are under-represented in the field error metric. As a consequence, the neural operator trained to minimize the field error naturally becomes agnostic to small scales, but in turbulent flows the small scales are important! On the contrary, NO+DM can capture both the large and small-scale turbulent structures. This indicates that the NO+DM model accurately captures the statistical structure of the turbulence compared to NO. Hence, higher field error and lower energy spectrum error indicate that the the underlying model respects the global statistical features and energy dynamics at the cost of some local phase alignment issues. Given this, we should also note that the field error of NO+DM is quite small in itself.
Field Error Energy Spectrum Error
NO NO+DM NO NO+DM
Kolmogorov Flow - FNO 5.94e-02 6.53e-02 9.09e-02 1.97e-03
Kolmogorov Flow - UNet 5.52e-02 6.04e-02 1.90e-02 2.57e-03
Kolmogorov Flow - TC-UNet 4.80e-02 5.38e-02 1.44e-02 1.99e-03
Buoyancy-driven Flow 1.20e-02 1.58e-02 8.10e-03 4.71e-04
Turbulent Airfoil Wake 8.35e-04 1.25e-03 8.82e-04 7.64e-04
Turbulent Jet 7.87e-03 1.57e-02 1.31e-02 1.02e-03
Table  1: Comparison of Field Error and Energy Spectrum Error for different methods and cases.
Diffusion model stabilizes autoregressive rollouts
Although neural operators are computationally efficient, performing long-term forecasts requires autoregressive rollouts where the output of the neural operator is fed back as the input. Unfortunately, error accumulates at each autoregressive step making longer rollouts unstable. The spectral bias of neural operators further aggreviates the quality of autoregressive rollouts. Here we investigate if diffusion models that mitigate spectral bias in neural operators have the ability to stabilize autoregressive rollouts. To this end, we train a UNet-based neural operator, [17], that takes a history of states - $\{u_{t-50\tau},u_{t-40\tau},...,u_{t}\}$ as the input and directly predicts the states - $\{u_{t+10\tau},u_{t+20\tau}\}$. We then auto-regressively predict the output upto $u_{t+300\tau}$. For the neural operator case, during the autoregressive (AR) rollout, we feed back the output predicted by the neural operator as its input to forecast further into the future. In the diffusion-corrected autoregressive (DCAR) rollout, the output of the neural operator is corrected by the diffusion model to recover the missing frequencies, before feeding it back as the input to the neural operator. We show b) the contour plots of two snapshots at $u_{t+120\tau}$ and $u_{t+200\tau}$ and c) the evolution of relative MSE across the auto-regressive steps in Fig 9. It can be seen that the NO (Neural Operator) prediction has some unrealistic artifacts, when predicting auto-regressively. The artifacts become more prominent as we forecast further into the future. On the other hand, the diffusion model helps stabilize longer rollouts into the future by reconstructing the missing frequencies at every autoregressive step. Moreover, it can be seen that initially the AR rollout predictions has lower Rel. MSE but as longer forecasts are carried out, at later time steps the Rel. MSE of the AR rollout exceeds that of the DCAR rollout. It is interesting to observe that at $\Delta t=120\tau$, the AR rollout with the unrealistic artifacts have a lower rel. MSE (2.25e-03) than DCAR rollout prediction (2.46e-03). The unexpected observation occurs because these high-frequency artifacts bear low energy and remain less significant while computing the relative MSE of the velocity field predictions. This behavior further amplifies the significance of the question - Is field MSE the perfect evaluation metric for turbulent systems.
[…]
@@
The energy spectra in Figure 2 to Figure 7 demonstrate that the diffusion model conditioned on the neural operator performs better than the neural operator alone. However, these plots are constructed for each time step separately and do not contain information about the cross-correlation of the states across time steps. We address this question by performing proper orthogonal decomposition (POD) on 10 timesteps predicted by the neural operator, as well as the diffusion model conditioned on the neural operator, and compare it with the corresponding state of the ground truth. Figure 10 illustrates the decay of the eigenvalues and compares the POD modes of the true data with the other two approaches for the Kolmogorov flow case. The proposed framework of the diffusion models conditioned on the neural operator has better aligned eigen spectrum and eigen modes with the ground truth than the neural operators alone. The POD analysis of buoyancy-driven flow (Figure A.1), turbulent airfoil (Figure A.2), 3D turbulent jet (Figure A.3) and Schlieren velocimetry of helium jet (Figure A.4) are also provided in Appendix A.
Fig.  10: POD analysis of Kolmogorov flow. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (FNO) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
@@
Summarizing, we propose a framework that utilizes neural operators as effective priors for diffusion models. Our results demonstrate that the proposed framework that blends neural operators with diffusion models significantly improves the energy spectrum compared to the neural operators alone and has been tested across different types of neural operator architectures - FNO, U-Net, and TC-UNet, over a variety of test cases. The framework utilized here is general, and therefore, any advances in the quality, computational efficiency, or scalability of either the neural operator or the diffusion models in the future can be applied seamlessly and could potentially improve the proposed framework.
@@ 4.1 Neural Operators
Neural Operators are a class of architectures in deep learning, that maps between infinite dimensional functional spaces. Consider an operator $\mathcal{G}$, that maps from the input function space $U$ to the output function space $V$, i.e., $\mathcal{G}:U\rightarrow V$. Then any machine learning architecture that is used to approximate this functional mapping to predict $v\in V$ for a given $u\in U$ can be defined as a neural operator. Neural operator is a popular framework for encoding dynamics and making predictions under some given conditions. But like any other neural network architectures, they suffer from spectral bias.
Lemma 4.1. Let $F:U\rightarrow V$ be a functional mapping with $U,V\in H^{s}(\Omega)$, where $H^{s}(\Omega)$ is the Sobelov space of functions defined on $\Omega$. Let $\mathcal{G}(\theta_{NO})$ be a neural operator parameterized by $\theta_{NO}$, that is approximating the functional space $F$. Then $\mathcal{G}$ learns the lower frequency components of the functional space $F$ more effectively [63, 64, 65, 66].
Sketch of Proof: We want to approximate the functional space $F(U)=V$, then if we take the Fourier transform for this function, we have
@@
where $\hat{F}(\xi)$ represents the amplitude of the the frequency component $\xi$. Now let $\mathcal{G}(\theta_{NO};U)$ is the operator approximating $F(U)$, typically the loss function $\mathcal{L}(\theta_{NO})$ used for optimizing such networks is given by
$\mathcal{L}(\theta_{NO})=\frac{1}{N}\sum_{n=1}^{N}(F(u_{n})-\mathcal{G}(\theta_{NO};u_{n}))^{2},$ (7)
where $u_{n}\in U$, and $n=1,2,\cdot\cdot\cdot,N$ are the discrete points where the functional evaluation are available for training. We can also express $\mathcal{G}$ in terms of its discrete Fourier components as
$\hat{\mathcal{G}}(\theta_{NO};\xi)=\frac{1}{N}\sum_{n=1}^{N-1}\mathcal{G}(\theta_{NO};u_{n})e^{-2\pi i\xi u_{n}/N},$ (8)
@@
$\mathcal{G}(\theta_{NO};u_{n})=\frac{1}{N}\sum_{n=1}^{N-1}\hat{\mathcal{G}}(\theta_{NO};\xi)e^{2\pi i\xi u_{n}/N}.$ (9)
@@
$\nabla_{\theta_{NO}}\mathcal{L}(\theta_{NO})=\frac{-2}{N}\sum_{n=1}^{N}(F(u_{n})-\mathcal{G}(\theta_{NO};u_{n}))\nabla_{\theta_{NO}}\mathcal{G}(\theta_{NO};u_{n}).$ (11)
@@
$\frac{d}{dm}(\hat{\mathcal{G}}(\theta_{NO};\xi))\propto-\alpha\hat{\nabla}_{\theta_{NO}}\mathcal{L}(\theta_{NO})\propto\frac{\partial\mathcal{G}(\theta_{NO};\xi)}{\partial\theta_{NO}}.$ (12)
Since typically in neural network architectures $\partial\mathcal{G}(\theta_{NO};\xi)/\theta_{NO}$ is smooth or piecewise smooth, $\alpha\hat{\nabla}_{\theta_{NO}}\mathcal{L}(\theta_{NO})$ is larger for smaller $\xi$ [63, 66, 67]. Therefore, the operator $\mathcal{G}$ learns the low frequency components of the functional space $F(U)$ more effectively compared to higher frequency components, which is known as the spectral bias. Hence like any other networks, operators suffer from spectral bias, which leads to ineffective refinement of the higher frequency component, leading, in turn, to sharp decay of the power spectrum associated with the output predicted by any neural operator. Although neural operators fail to capture the high frequency, it captures effectively the mean flow dominated by the low wave numbers. This acts as an ideal prior for diffusion models to mimic the true flow consisting of all the frequencies.
@@ 4.2 Score-based diffusion models
Diffusion models [38] are generative algorithms that can effectively generate samples from the true underlying distribution of functions, $\mathcal{T}$, given a finite set of samples from $\mathcal{T}$. The diffusion models generate samples from $\mathcal{T}$ by learning to travel from a simple distribution such as a standard normal distribution ($\Gamma_{0}=\mathcal{N}(0,I)$) to the desired complex distribution ($\Gamma_{N}\approx\mathcal{T}$) through $N$ iterative steps. Specifically, at first the diffusion model takes the function $\mathbf{X_{0}}$ sampled from $\Gamma_{0}$ as the input and predicts $\mathbf{X}_{1}$. Since $\Gamma_{0}=\mathcal{N}(0,I)$, $\mathbf{X}_{0}$ can be easily sampled. Now, the diffusion model can estimate $\mathbf{X}_{i+1}$ from $\mathbf{X}_{i}$ iteratively for $N$ steps. But how can we train the diffusion model to iteratively transition from $\Gamma_{0}=\mathcal{N}(0,I)$ to $\Gamma_{N}\approx\mathcal{T}$, the desired true distribution, when we do not have ground truth samples from the intermediate distributions $\Gamma_{i}$ for $i=\{1,2,3,\cdots,n-1\}$? This can be achieved through denoising score matching with Langevin dynamics [39].
@@
The overall goal of a score ($s$) based diffusion model is to estimate the score function defined by $s_{\theta_{D}}(\mathbf{X})=\nabla_{X}\log p(\mathbf{X})$, where $\theta_{D}$ are the parameters of the diffusion model and $p$ is the probability density of $\mathbf{X}$, where $\mathbf{X}$ is the continuous representation of $\mathbf{X}_{i}\sim\Gamma_{i}$. Since the data distribution is not known explicitly, and there is a chance that the data might lie on a lower dimensional manifold, the score function may be ill defined in regions where there is no data. To circumvent this issue, perturbation of the data is done by adding Gaussian noise, which ensures that the score is well defined across the entire space by smoothing the distribution. Hence, the score function gives us a direction to move towards the region of higher probability. However the direct way to sample from the distribution is missing here. In order to fix this problem, Langevin Dynamics was used in [68]. Langevin Dynamics ensures the convergence of the generated samples to the true underlying distribution. It balances the inherent dynamics between deterministic movement, which is driven by the gradient of log of the probability, with random exploration given by the noise levels. In our proposed method, we use the score function, which is additionally conditioned on the output of the neural operator $\mathbf{\mathcal{G}}(\theta_{NO};(\mathbf{x,t}))$ (for readability expressed as $\mathcal{G}$) given by
$s_{\theta_{D}}(\mathbf{X},\sigma,\mathbf{\mathcal{G}})=\nabla_{X}\log p(\mathbf{X}|\mathbf{\mathcal{G}}),$ (13)
@@
where $\sigma$ is the noise level. The modified score function now directs the diffusion model to sample from the posterior distribution of $\mathbf{X}$ given $\mathbf{\mathcal{G}}$. This ensures that the the generated samples from the diffusion model are consistent with both the structures defined by $\mathbf{\mathcal{G}}$ and the data distribution that is learned by the diffusion model. The update rule for Langevin dynamics can be shown as
$\mathbf{X}_{i+1}=\mathbf{X}_{i}+\frac{\varepsilon}{2}s_{\theta_{D}}(\mathbf{X}_{i},\sigma_{j},\mathbf{\mathcal{G}})+\sqrt{\varepsilon}z_{i},$ (14)
@@
where $\varepsilon$ is the step size, $z_{i}$ is the noise component and $\sigma_{j}$ is the scale of noise at the particular time step $j$ during the diffusion process. From Eqns 13 and 14, we can see that as $\sigma_{i}$ decreases, the score function $s_{\theta_{D}}$ focuses on recovering the high-frequency details which was smoothen out by the noise. In the frequency domain Eqn 14 can be shown as
$\hat{\mathbf{X}}_{i+1}=\hat{\mathbf{X}}_{i}+\frac{\varepsilon}{2}\hat{s}_{\theta_{D}}(\mathbf{X}_{i},\sigma_{j},\mathbf{\mathcal{G}})+\sqrt{\varepsilon}\hat{z}_{i},$ (15)
@@
$u(\mathbf{x,t})=\sum_{n=1}^{\infty}a_{n}(\mathbf{t})\phi_{n}(\mathbf{x}),$ (16)
where $\phi_{n}(\mathbf{x})$ are the basis functions and $a_{n}(\mathbf{t})$ are the time-dependent coefficients.
Let $\mathbf{\mathcal{G}}$ and $s_{\theta_{D}}(\mathbf{X},\mathbf{\mathcal{G}},\sigma)$ represents the output of the neural operator and the diffusion model respectively, such that
$s_{\theta_{D}}(\mathbf{X},\mathbf{\mathcal{G}},\sigma)\simeq\sum_{n=1}^{\infty}a_{n}^{s}(\mathbf{t})\phi_{n}(\mathbf{x}).$ (17)
The spectral bias suffered by neural operators limits its ability to learn the true solutions beyond a critical wavenumber $|k_{c}|$. Let $N_{c}$ represent the critical basis index in the spectrally decomposed representation of $\mathcal{G}$ beyond which the neural operator’s prediction carries trivial amount of energy. In this context, $|k_{c}|$ can interpreted as the wavenumber that carries maximum energy in the true mode $\hat{\phi}_{N_{c}}(k)$. The energy of $\mathcal{G}$, $E_{\mathcal{G}}(t,k)$ can now be represented as,
$E_{\mathcal{G}}(t,k)=\frac{1}{N}\left|\sum_{n=1}^{N}a_{n}^{\mathcal{G}}(t)\hat{\phi}_{n}(k)\right|^{2}\approx\frac{1}{N}\left|\sum_{n=1}^{N_{c}}a_{n}^{\mathcal{G}}(t)\hat{\phi}_{n}(k)\right|^{2},$ (18)
because $\left|\sum_{n=N_{c}+1}^{N}a_{n}^{\mathcal{G}}(t)\hat{\phi}_{n}(k)\right|^{2}\approx 0$. From Figure 2 we see that using a better neural operator has a better aligned energy spectrum implying a larger $k_{c}$. Subsequently, the output of the diffusion model also improves as shown in Table 1. In other words, a better $\mathcal{G}$ serves as an improved prior increasing $k_{c}$ and $\left|u(\mathbf{x},t)-s_{\theta_{D}}(\mathbf{x},\mathbf{\mathcal{G}},\sigma))\right|\rightarrow 0$.
@@
$s_{\theta_{D}}(\mathbf{X},\mathbf{\mathcal{G}},\sigma)$ $\simeq\sum_{n=1}^{\infty}a_{n}^{s}(t)\phi_{n}(\mathbf{X})$ (19)
$=\sum_{n:|k|\leq|k_{c}|}a_{n}^{s}(t)\phi_{n}(\mathbf{x})+\sum_{n:|k|>|k_{c}|}a_{n}^{s}(t)\phi_{n}(\mathbf{x}).$ (20)
Let us assume, in the discrete Fourier space we can write it in terms of the energy spectrum as
$E_{s_{\theta_{D}}}(k)=\frac{1}{N}\left|\sum_{(n\leq N_{c})}a_{n}^{s}(t)\hat{\phi_{n}}(k)\right|^{2}+\frac{1}{N}\left|\sum_{(n>N_{c})}a_{n}^{s}(t)\hat{\phi_{n}}(k)\right|^{2},$ (21)